Given closed economy model: ( )

i. Drive the balanced budget multiples and show it does not have multiples effect.

(Hint dG = dT)

ii. Find the government multiplier for a variable tax

iii. Suppose the lump sum tax rate has increased. Derive the tax rate multiplier and show

increase in tax rate is regressive.


1
Expert's answer
2022-01-25T08:32:08-0500

A)The tax multiplier measures the impact of a tax reform on aggregate demand. Reduced taxes have the same effect on income and consumption as increased government spending.


The formula for the simple spending multiplier is :"\frac{1}{mps}"


"B)Tax \ Multiplier = \frac{ΔY }{ΔT} = \frac{– MPC }{1 – MPC}"


"C)increase\ in\ Tax\ Receipt\ (ΔT) =\frac{ ΔY }{ Tax Multiplier}"